OpenStudy (anonymous):

Select all of the following expressions that are equal to 2.5. 5 ÷ 2 -2(-1.25) (-0.5)2 |-1 + (-1.5)|

OpenStudy (muscrat123):

5 / 2 = 2.5

OpenStudy (anonymous):

thats a right?

OpenStudy (muscrat123):

-2 x -1.25 = 2.50 and yes

OpenStudy (anonymous):

okay so a and b

OpenStudy (muscrat123):

-0.5 x 2 = 1.0

OpenStudy (anonymous):

not c

OpenStudy (muscrat123):

and can u solve d? i dont want to give u all the answers ps: it = to 2.5-i want u to find out how i got there

OpenStudy (anonymous):

negatives are canceled right?

OpenStudy (muscrat123):

do u know what \[\left| these \right|\] mean?

OpenStudy (anonymous):

its means it cancels the negatives right

OpenStudy (muscrat123):

not quite - they mean the absolute value of something. do u know what an absolute value is?

OpenStudy (anonymous):

like the first 1 right? not the 1 in 1.5

OpenStudy (muscrat123):

not quite right either - an absolute value is the distance away from 0. so if an ENTIRE problem is in \[\left| these \right|\],then it doesnt matter if it is negative or positive. it will be positive ALWAYS if the ENTIRE problem is in the symbol do u understand do far?

OpenStudy (anonymous):

yea

OpenStudy (muscrat123):

ok. so solve for inside the symbols. in this case the ENTIRE problem r in the symbols, so u can ignore the - and positive signs i put emphasis on "entire" bc if the ENTIRE problem is not in the symbols, the rules r different

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so d is one bacause that equaks 2.5

OpenStudy (anonymous):

*equals

OpenStudy (muscrat123):

yes

OpenStudy (muscrat123):

anymore?

OpenStudy (anonymous):

yup

OpenStudy (muscrat123):

how many?

OpenStudy (anonymous):

i have no clue over like 200

OpenStudy (muscrat123):

*mouthdropped

OpenStudy (muscrat123):

for what class?

OpenStudy (anonymous):

i have a whole semester too make up and thats not including science

OpenStudy (muscrat123):

whole semester from what?

OpenStudy (anonymous):

algebra 1 and physics

OpenStudy (muscrat123):

and i may b able to help w science

OpenStudy (anonymous):

the whole 6 weeks thing

OpenStudy (muscrat123):

what school?

OpenStudy (muscrat123):

*mouth dropped in unbelievement

OpenStudy (anonymous):

regulare high school

OpenStudy (muscrat123):

?

OpenStudy (muscrat123):

online?

OpenStudy (anonymous):

no just plain old high schooll but i am makeing up online

OpenStudy (muscrat123):

oh. how did u fall behind?

OpenStudy (anonymous):

its complicated

OpenStudy (muscrat123):

lol ok

OpenStudy (muscrat123):

srry 4 drillin u

OpenStudy (muscrat123):

next ?

OpenStudy (anonymous):

im getting it

OpenStudy (anonymous):

i just dont like talking about it

OpenStudy (muscrat123):

oh ok. srry. and when i said next ? , i meant next question. i just abbreviate things

OpenStudy (anonymous):

oh yea im getting it i just have to type out the fractions

OpenStudy (muscrat123):

cant u just copy and paste?

OpenStudy (anonymous):

yea but it doesnt take fractions with it

OpenStudy (muscrat123):

it just puts it in a/b form